Two forces acting on an object have magnitudes F1 6.6 N and F2 2.2 N. Which third force causes the object to be in equilibrium? F2 a) 4.4 Na 1620 counterclockwise from F b) 4.4 Na 1080 counterclockwise from F c) 7 N 162° counterclockwise from F d) 7 N 108o counterclockwise from Fi
Answer #1

here,

F1 = 6.6 N (- i)

F2 = 2.2 N ( j)

let the third force required be F3

as they are in equilibrium

F1 + F2 + F3 = 0

- 6.6 N i + 2.2 N j + F3 = 0

F3 = 6.6 N i - 2.2 N j

the magnitude of force , |F3| = sqrt(6.6^2 + 2.2^2) = 7 N

the direction of force , theta = arctan((-2.2)/6.6) = 162 degree counterclockwise from F1

so the correct option is &)